# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 2.1 Structural Result for Multiclass Linear Classification
$\sum_{i:\mathbf{w}^{(i)}\neq\mathbf{0}}\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\;.$ (2)
Moreover, if for some $i\in[K]$, we have that $\gamma-t_{i}\leq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ then $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{i}(\mathbf{x})\geq\gamma]=\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq(\gamma-t_{i})/\left\|\mathbf{w}^{(i)}\right\|]\geq\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq\epsilon]$. However, for $\epsilon<2\sqrt{\log({48})}$ from A.1 we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]>1/{48}$ which leads to a contradiction. Hence, we have that $\gamma-t_{i}\geq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ for all $i\in[K]$. Now from the fact that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\gtrsim C\cdot\mathrm{OPT}+\epsilon\gtrsim\epsilon$.
@@
Concluding, from (2) we have that there exists $i\in[K]$ such that $\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon$. Now since $\gamma-t_{i}\geq 0$, we have that $\mathbf{w}^{(i)}\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\geq 0$ and hence $\left(\mathbf{w}^{(i)}/\left\|\mathbf{w}^{(i)}\right\|\right)\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon^{2}$. Using the fact that we centered the affine functions defining $f$ by subtracting $\mathbf{w}^{(1)}\cdot\mathbf{x}+t_{1}$ from each of them, gives us the statement. This completes the proof of Proposition 2.4.
@@ 2.2 Finding a Relevant Direction
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{2}/({K^{2}\sqrt{k}})$, for $c$ a sufficiently small constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$. Let $f$ be a multiclass linear classifier with weight vectors $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $\mathrm{OPT}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=d\mathrm{poly}(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ and a universal constant $C$ such that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$, then Algorithm 2 with probability at least $1-\delta$ returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(1/\epsilon,K)$, such that: There exists $\mathbf{v}\in\mathcal{E}$ and $i,j\in[K],i\neq j$ for which
@@
Before proving Proposition 2.8, we develop the necessary tools for working with multiclass linear classifiers restricted to cubes extending along the orthogonal complement of a subspace. First, we define the notion of a projected classifier onto a set of linear equations, Definition 2.9.
@@
Let $f(\mathbf{x})=\operatorname*{argmax}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$ be a multiclass linear classifier, let $\mathcal{G}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ be a set of $k$ linear equations and let $V$ be the subspace spanned by orthonormal vectors $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\in\mathbb{R}^{d}$. We denote by $f_{\mathcal{G}}(\mathbf{x})$ the linear classifier $f$ projected onto $\mathcal{G}$, i.e., $f_{\mathcal{G}}(\mathbf{x})=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp_{V}}\cdot\mathbf{x}+t_{i}+t_{i,\mathcal{G}}^{\prime})$, where $t_{i,\mathcal{G}}^{\prime}=\sum_{j=1}^{k}z_{j}a_{j}$ where $a_{j}=\mathbf{w}^{(i)}\cdot\mathbf{v}^{(j)}$.
@@
Let $d,K\in\mathbb{Z}_{+}{,\epsilon>0}$, $f:\mathbb{R}^{d}\to[K]$ be a multiclass linear classifier, $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d},{k\geq 1}$ and $\epsilon>0$. Let $\mathcal{S}$ an $\epsilon$-approximating partition of $\mathbb{R}^{d}$ with respect to $V$ as in Definition 2.2. Then there exists a set $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq 1-\epsilon$ and for any $S\in\mathcal{T}$, it holds that
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\leq K^{2}{\sqrt{k}}\epsilon\;,$
@@
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\dots,t_{K}\in\mathbb{R}$ such as $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. Denote by $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$ an orthonormal basis of $V$. Fix a set of the partition $S\in\mathcal{S}$ and denote by $f_{\mathcal{G}_{S}}=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp{V}}\cdot\mathbf{x}+t_{i}+t_{i}^{\prime})$ the projected classifier that corresponds to the set of linear equations $\mathcal{G}_{S}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ (see Definition 2.2). Let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, $t_{i,j}=t_{i}-t_{j}$ and $t_{i,j}^{\prime}=t_{i}^{\prime}-t_{j}^{\prime}$. By the union bound, we have that
@@
Furthermore, by the definition of $f_{\mathcal{G}_{S}}$, if $\mathbf{x}\in S$ then we have that for any $i,j\in[K]$
$\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$ $=\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|$
$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\lvert\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\rvert$
$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\epsilon$
$\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\;,$
where we used the fact that the vector $\sum_{i=1}^{k}z_{i}\mathbf{v}^{(i)}$ belongs in $S$ by the definition of $\mathcal{G}$.
@@
Notice that in order for $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$, we need to have that $\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$, because we can rewrite $t^{\prime}_{i,j}=\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}+t_{i,j}^{\prime}$.
Averaging over $S\in\mathcal{S}$ gives us
@@
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$ $\leq\sum_{i\neq j\in[K]}\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$
@@
$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon]$
$\leq\sum_{i\neq j\in[K]}\frac{\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon}{\|\mathbf{w}^{(i,j)}\|}\leq K^{2}{\sqrt{k}}\epsilon\;.$
@@
Let $\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\ldots,t_{K}\in\mathbb{R}$ be the parameters defining $f$, i.e., $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. Denote by $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, by $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}=(\mathbf{w}^{(i,j)})^{\perp V}/\left\|(\mathbf{w}^{(i,j)})^{\perp V}\right\|$, and let $C$ be a constant that satisfies the statement of Proposition 2.4. Let $\mathcal{G}_{S}$ be the set of linear equations associated with each cube $S\in\mathcal{S}$ as defined in Definition 2.2, and let $f_{\mathcal{G}_{S}}$ be the projection of $f$ onto $\mathcal{G}_{S}$ as defined in Definition 2.9. Furthermore, denote by $\mathbf{u}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ the population moments of each discretization set and by $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ its empirical counterpart, computed in Line 4. Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 5 of Algorithm 2.
We first show that if the current subspace $V$ does not lead to a good approximation of $f$, then by splitting the space into small cubes with respect to $V$, there will be a nontrivial fraction of them for which $f$ is far from constant.

### v2

@@ 2.1 Structural Result for Multiclass Linear Classification
$\sum_{i:\mathbf{w}^{(i)}\neq\mathbf{0}}\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\gtrsim\epsilon\;.$
This implies that there exists some $j\in[K]$ such that
[…]
To conclude the proof, we must show that $\gamma-t_{j}\geq\epsilon\,\left\|\mathbf{w}^{(j)}\right\|$. Suppose, for the sake of contradiction, that for some $i\in[K]$ we have $\gamma-t_{i}\leq\epsilon\left\|\mathbf{w}^{(i)}\right\|$. Then, by the definition of $L_{i}$, we have
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}L_{i}(\mathbf{x})\geq\gamma\Big{]}=\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}\Bigg{[}x\geq\frac{\gamma-t_{i}}{\left\|\mathbf{w}^{(i)}\right\|}\Bigg{]}\geq\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq\epsilon]\;.$
However, for $\epsilon<2\sqrt{\log(48)}$, A.1 implies that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma\Big{]}>\frac{1}{48},$
which leads to a contradiction from Lemma 2.7 (and the choice of $\gamma$). Thus, we must have $\gamma-t_{i}\geq\epsilon\,\left\|\mathbf{w}^{(i)}\right\|$, for all $i\in[K]$. As a result, we obtain
[…]
@@
Finally, using the fact that we centered the affine functions defining $f$ by subtracting $\mathbf{w}^{(1)}\cdot\mathbf{x}+t_{1}$ from each of them, we obtain the desired statement. This completes the proof of Proposition 2.4.
@@ 2.2 Finding a Relevant Direction
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k})$, for some sufficiently small absolute constant $c>0$. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$. Let $f$ be a multiclass linear classifier with weight vectors $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and define $\mathrm{OPT}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. Furthermore, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ defined relative to the subspace $V$ (as in Definitions 2.2 and 2.3).
There exists a sample size $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ and a universal constant $C>0$, such that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$, then Algorithm 2 given $N$ samples, runs in $\mathrm{poly}(N)$ time and with probability at least $1-\delta$ returns a set of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(1/\epsilon,K)$ with the following property: there exists a $\mathbf{v}\in\mathcal{E}$ and $i,j\in[K]$, $i\neq j$, such that
@@
Before proving Proposition 2.8, we develop the necessary tools for working with multiclass linear classifiers restricted to cubes extending along the orthogonal complement of a subspace. First, we define the notion of a projected classifier onto a set of linear equations (Definition 2.9).
@@
Let $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$, let $\mathcal{G}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots,\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ be a set of $k$ linear equations and let $V$ be the subspace spanned by orthonormal vectors $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\in\mathbb{R}^{d}$. We denote by $f_{\mathcal{G}}(\mathbf{x})$ the linear classifier $f$ projected onto $\mathcal{G}$, i.e., $f_{\mathcal{G}}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp_{V}}\cdot\mathbf{x}+t_{i}+t_{i,\mathcal{G}}^{\prime})$, where $t_{i,\mathcal{G}}^{\prime}=\sum_{j=1}^{k}z_{j}a_{ij}$ where $a_{ij}=\mathbf{w}^{(i)}\cdot\mathbf{v}^{(j)}$.
@@
Let $d,K,k\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}>0$, $f:\mathbb{R}^{d}\to[K]$ be a multiclass linear classifier, $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d},k\geq 1$. Let $\mathcal{S}$ be an $\epsilon^{\prime}$-approximating partition of $\mathbb{R}^{d}$ with respect to $V$ as in Definition 2.2, with $\epsilon^{\prime}\leq\epsilon^{2}/(K^{2}\sqrt{k})$. Then there exists a set $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq 1-\epsilon$ and for any $S\in\mathcal{T}$, it holds that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\leq\epsilon\;,$
@@
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\dots,t_{K}\in\mathbb{R}$ such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. Let $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$ be an orthonormal basis of $V$. Fix a set of the partition $S\in\mathcal{S}$ and denote by $f_{\mathcal{G}_{S}}=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp{V}}\cdot\mathbf{x}+t_{i}+t_{i}^{\prime})$ (see Definition 2.9 for the definition of $t_{i}^{\prime}$) the projected classifier that corresponds to the set of linear equations $\mathcal{G}_{S}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots,\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ (see Definition 2.2 for the definition of $\mathcal{G}_{S}$). Let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, $t_{i,j}=t_{i}-t_{j}$ and $t_{i,j}^{\prime}=t_{i}^{\prime}-t_{j}^{\prime}$. By the union bound, we have that
@@
Fix $i,j\in[K]$ with $i\neq j$. Notice that for a sign change to occur, i.e., $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$ it is necessary that
$\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\Bigr{|}\leq\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\Bigr{|}.$
To see that, note that we can add and subtract $\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}$, and we get that if for some $\mathbf{x}\in S$ we have $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$ we also have, $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j}-\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V})$, hence it should hold that $\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\Bigr{|}\leq\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\Bigr{|}$. Therefore, it suffices to bound the probability that $\operatorname*{\mathbf{Pr}}[|\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}|\leq|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}|\mid\mathbf{x}\in S]$. By the definition of $f_{\mathcal{G}_{S}}$, if $\mathbf{x}\in S$ then for any $i,j\in[K]$ we have
$\left|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\right|=\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|.$
This equality holds because by definition the projection of $\mathbf{x}$ onto the subspace $V$ with orthonormal basis $\{\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\}$ is given by $\mathbf{x}^{V}=\sum_{l=1}^{k}(\mathbf{x}\cdot\mathbf{v}^{(l)})\mathbf{v}^{(l)}$, and the term $t_{i,j}^{\prime}$ is defined as $t_{i,j}^{\prime}=\sum_{l=1}^{k}z_{l}\,(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})$. Next, by applying the triangle inequality we have
$\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|\leq\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\Bigl{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Bigr{|}.$
By the definition of the partition $S$, if $\mathbf{x}\in S$ then for every $l\in\{1,\ldots,k\}$ the projection of $\mathbf{x}$ onto the direction $\mathbf{v}^{(l)}$ is $\epsilon^{\prime}$-close to the corresponding value $z_{l}$, i.e., $\Big{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Big{|}\leq\epsilon^{\prime}$. Therefore,
$\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\Bigl{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Bigr{|}\leq\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\epsilon^{\prime}.$
Finally, by applying the Cauchy-Schwarz inequality, we obtain
$\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\leq\sqrt{k}\,\sqrt{\sum_{l=1}^{k}\Bigl{(}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{)}^{2}}=\sqrt{k}\,\|(\mathbf{w}^{(i,j)})^{V}\|\;.$
Combining the above, we have $\left|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\right|\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\,\epsilon^{\prime}$. Averaging over $S\in\mathcal{S}$ gives us
@@
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$ $\leq\sum_{i\neq j\in[K]}\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\epsilon^{\prime}\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$
@@
$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\epsilon^{\prime}]$
$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|\mathbf{w}^{(i,j)}\|\sqrt{k}\epsilon^{\prime}]$
$\leq\sum_{i\neq j\in[K]}\sqrt{k}\epsilon^{\prime}\leq K^{2}\sqrt{k}\epsilon^{\prime}\;,$
where we used the anti-concentration of the Gaussian distribution (see A.1). Finally, since $\epsilon^{\prime}\leq\epsilon^{2}/(K^{2}\sqrt{k})$ it holds that $\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\leq\epsilon^{2}$. To conclude the proof, let ${\mathcal{S}}_{bad}$ be the set of $S\in\mathcal{S}$ such that if $S\in{\mathcal{S}}_{bad}$ then $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]>\epsilon$. Then, we have that
$\epsilon^{2}\geq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\geq\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]>\epsilon\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[S]\;.$
Hence, $\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[S]\leq\epsilon$. By taking $\mathcal{T}=\mathcal{S}/{\mathcal{S}}_{bad}$ we conclude the proof. ∎
@@
Let $\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\ldots,t_{K}\in\mathbb{R}$ be the parameters defining $f$, i.e., $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. For brevity, define the difference vectors $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, and let $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}=(\mathbf{w}^{(i,j)})^{\perp V}/\left\|(\mathbf{w}^{(i,j)})^{\perp V}\right\|$ denote the normalized projection of $\mathbf{w}^{(i,j)}$ onto the orthogonal complement of a given subspace $V$. Also, define $W=\mathrm{span}(\mathbf{w}_{1},\dots,\mathbf{w}_{K})$, and let $C$ be a constant as specified in Proposition 2.4. For each cube $S\in\mathcal{S}$ (as defined in Definition 2.2), let $\mathcal{G}_{S}$ be the corresponding set of linear equations, and let $f_{\mathcal{G}_{S}}$ be the projection of $f$ onto $\mathcal{G}_{S}$ (see Definition 2.9). Furthermore, for every $S\in\mathcal{S}$ and $i\in[K]$, define the population moment $\mathbf{u}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$,and let $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$ its empirical counterpart (see Line 4). Also, denote by $\widehat{\mathbf{U}}$ the matrix computed at Line 5 of Algorithm 2.
We first show that if the current subspace $V$ does not lead to a good approximation of $f$, meaning that any function defined only on $V$ gets error at least $\Omega(\mathrm{OPT}+\epsilon)$ with the labels $y$, then when we partition the space into small cubes with respect to $V$, on a nontrivial fraction of them $f$ is far from being constant.